OpenStudy (anonymous):

@jeweaver6895 This is the last thing I need help with. I promise this is the last thing I need, scouts honor. Thanks a billion! :]

OpenStudy (anonymous):

Its no problem, im happy to help :) ok this one your going to use your trig identities again. but lets begin with the angle, what do you know about the sum of the angles in a triangle?

OpenStudy (anonymous):

They all equal to 180 degrees.

OpenStudy (anonymous):

correct, so when one angle is 50 and one is 90, what does the third angle have to be?

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

Right! ok now onto the sides! so lets work with angle A as we try to find the values of the sides of the triangle

OpenStudy (anonymous):

For side AB we know that it is the hypotenuse, and we have a side adjacent to angle A, so which trig identity should we use?

OpenStudy (anonymous):

Cos? adjacent/hypotenuse

OpenStudy (anonymous):

Correct! so set that up for me, how do you use it?

OpenStudy (anonymous):

cos 5/125 sorry if its wrong it honestly my best guess

OpenStudy (anonymous):

dont be sorry if its wrong, failures are how we learn :) So we know that the cos(50)=125/AB because cos(x)=adj/hyp

OpenStudy (anonymous):

Ohh okay I see how it done now makes sense. :] Also thanks sorry it's just I always feel like apologizing if I make a mistake.

OpenStudy (anonymous):

Great so what did you get for AB? its alright :)

OpenStudy (anonymous):

I got 66, is that right or wrong?

OpenStudy (anonymous):

no thats not quite it :) so did you divide 125 by cos(50) to solve for AB?

OpenStudy (anonymous):

ohhhhhhh okay give me a sec I see what I did wrong

OpenStudy (anonymous):

194?

OpenStudy (anonymous):

YEP! ok now the last one! how do you find CB?

OpenStudy (anonymous):

I use the pythagorean theorem now?

OpenStudy (anonymous):

you could use that, it might be a little easier to use tan

OpenStudy (anonymous):

Well I got 231 using the theorem was it right or wrong?

OpenStudy (anonymous):

no thats not right, CB has to be less than the hypotenuse, try the other method :)

OpenStudy (anonymous):

I got 33 but it doesn't look right to me so I think I'm doing this wrong but I'm not exactly sure what could you point out what it is if you can tell?

OpenStudy (anonymous):

no thats not quite it either, ok ill walk you through it :) so from trig we know that the tan(50) is the opposite/adjacent, so when we put in values we get tan(50)=CB/125

OpenStudy (anonymous):

rearranging the equation a little bit we find that CB=125*tan(50)

OpenStudy (anonymous):

Ohhhhhhhhhhh I was dividing the whole time and its actually multiplying so now I got 148.969 which rounds to 149.

OpenStudy (anonymous):

ahh thatll get ya! you got it! Congratulations! :)

OpenStudy (anonymous):

Thanks so much again!

OpenStudy (anonymous):

Yeah you bethca!